Consider the closed economy with C = 800, G = 250, S = 100, and the private saving = -50 (notice “-“).

2.a. Answer the total income and the investment.

Total income:

Investment:



2.b. Answer the tax revenue and the public saving. Does the public saving increase or decrease the national saving?

Tax revenue:

Public saving:

Increase or Decrease?
1
Expert's answer
2020-04-16T09:22:18-0400

Total income

"Y=C+G+S"

"=800+250+100-150"

"=1000"


Investment

"I=Y-G-C"

"=1000-800-250"

"=-50"

Public saving decreases the national saving